# Machine-learning approaches to dispersion measure estimation for fast radio bursts

### v1

@@ Abstract
pubyear: 2025pagerange: Machine-learning approaches to dispersion measure estimation for fast radio burstsD
@@ 1 Introduction
In this paper, we investigate the use of machine-learning (ML) techniques to improve dedispersion and DM estimation for FRBs. While convolutional neural networks (CNNs; Connor & van Leeuwen 2018) have previously been proposed as substitutes for traditional dedispersion backends, most ML studies to date have focused on the classification or detection of FRBs rather than on quantitative DM estimation. Here, we extend this exploration by developing and comparing three deep-learning architectures, a baseline CNN, a fine-tuned ResNet-50, and a hybrid CNN–LSTM model, designed to recover the intrinsic DM directly from frequency–time dynamic spectra. Section 2 reviews previous ML applications to FRBs, Section 3 describes the generation of synthetic FRB data used for training and validation, and Section 4 outlines the architecture and training procedures for the three models. Our results are presented and discussed in Section 5, and the main conclusions and future directions are summarised in Section 6. Additional details on the simulated data, cross-validation, and test-set performance are provided in the Appendices at the end.
@@ 2 FRBs and Machine Learning
As previously stated, ML techniques have increasingly been applied to FRBs, though most studies have focused on classification rather than quantitative parameter estimation. Existing applications fall into two main categories: distinguishing genuine FRB signals from radio-frequency interference (RFI) or known pulsars (Wagstaff et al., 2016; Akeret et al., 2017; Mosiane et al., 2017; Czech et al., 2018a; Agarwal et al., 2020), and differentiating between repeating and non-repeating sources (Yang et al., 2023; Zhu-Ge et al., 2023; Luo et al., 2023; García et al., 2024; Sharma & Rajpaul, 2024; Qiang et al., 2025; Saliwanchik & Slosar, 2022). Various ML models have been explored in these contexts. For example, Wagstaff et al. (2016) used a Random Forest classifier to enhance FRB detection within the VLBA’s V-FASTR project (Wayth et al., 2011), while Akeret et al. (2017) applied CNNs to discriminate genuine FRBs from RFI and pulsars. Saliwanchik & Slosar (2022) further used neural networks to classify FRBs as repeating or non-repeating.
A notable example is the work of Agarwal et al. (2020), who demonstrated the use of CNNs and transfer learning to improve FRB detection. They fine-tuned pre-trained image-classification networks such as VGG16 (Simonyan & Zisserman, 2015), ResNet50 (He et al., 2016a), and Xception (Chollet, 2017) on a combined dataset of simulated FRBs, real RFI events from the Green Bank Observatory, and pulsar data. Their approach incorporated both frequency–time and DM–time representations to capture spectro-temporal and dispersion-related features, introducing a fusion technique that combined information from the two input domains to enhance classification accuracy.
Interest in ML-driven FRB analysis continues to grow, but the development of more sophisticated models has been limited by the scarcity of publicly available, high-quality data. To overcome this, many studies have generated synthetic data to supplement real observations. For instance, Connor & van Leeuwen (2018) developed a deep CNN trained on dynamic spectra, pulse profiles, and DM–time data using simulated FRBs injected into actual survey noise. They also explored whether CNNs could eventually replace traditional real-time dedispersion backends, but emphasized that low signal-to-noise per pixel remains a fundamental limitation for such applications with current telescopes.
@@ 3.1 Data Simulation
Scattering effects, which both broaden the FRB pulse and slightly delay its peak, were incorporated by convolving the intrinsic Gaussian profile with a one-sided exponential kernel,
@@
Figure 1: Example of a simulated FRB dynamic spectrum with $\mathrm{DM}\approx 559.388~\mathrm{pc\,cm^{-3}}$.
@@ 4.1 Baseline Convolutional Neural Network (CNN) Model
The baseline architecture follows the original design of a convolutional neural network (CNN; LeCun et al. 1998) developed by Rajabi (2024), implemented here for DM estimation from frequency–time dynamic spectra. The model comprises seven convolutional layers followed by a sequence of fully connected (dense) layers, as summarised in Table 1. Each convolutional layer applies a set of filters (kernels) to detect local spectro–temporal features and hierarchically capture complex patterns in the input data. To stabilise training and accelerate convergence, batch normalisation (Ioffe & Szegedy, 2015) is applied after each convolutional layer, followed by a Rectified Linear Unit (ReLU) activation (Nair & Hinton, 2010) to introduce non-linearity. Max-pooling operations with $2\times 2$ kernels are used between most convolutional layers (except the third and sixth) to reduce the spatial resolution of feature maps while retaining the most significant features, thereby improving computational efficiency and generalisation.
The convolutional layers progressively increase in depth, starting with 32 filters in the first layer and doubling in number through successive layers, culminating in 2048 filters in the seventh (final) convolutional layer, following the same principle as VGGNet (Visual Geometry Group; Simonyan & Zisserman 2014). This hierarchical structure enables the network to learn increasingly abstract representations, from simple spectral features in early layers to complex temporal correlations in deeper ones. Following the convolutional block, a series of fully connected (FC) layers integrates the extracted features to produce a single scalar output corresponding to the predicted DM. The dense section reads the flattened activations (100,352 elements) from the convolutional block, encodes them into 16,384 outputs, and then halves in size through successive layers until reaching the output neuron. The complete layer-by-layer configuration is summarised in Table 1 (adapted from Rajabi 2024), where the numbers of inputs and outputs ($a{\to}b$) indicate the number of feature maps entering and leaving each convolutional or dense layer.
@@
During hyper-parameter tuning, we explored architectures ranging from four to seven convolutional layers and varied the widths of the dense layers to achieve an optimal balance between model complexity and overfitting. Batch normalisation significantly improved convergence stability by reducing internal covariate shift (Ioffe & Szegedy, 2015), enabling the use of higher learning rates without divergence. Although dropout is a common regularisation technique (Srivastava et al., 2014), tests with dropout probabilities between 0 and 0.2 showed that random unit removal degraded performance, since strong spatial dependencies exist between neighbouring pixels in the input data (Zhang et al., 2018). Consequently, dropout layers were omitted.
@@
The model minimises the MAE loss function, which is well suited to continuous regression tasks such as DM estimation. We used the Adam algorithm (Adam et al., 2014) as the optimiser. Adam adapts learning rates for individual parameters based on mini-batch updates, making it well suited for large datasets in deep learning. Learning rates in the range $10^{-7},10^{-2}]$ were tested, with the smallest value ($10^{-7}$) yielding the smoothest convergence and most accurate results. Larger learning rates led to divergence from the optimal solution, whereas smaller rates allowed for gradual and stable parameter updates.
@@ 4.2 ResNet-50
ResNet-50, a deep convolutional neural network comprising 50 layers (He et al., 2016b), was fine-tuned to adapt its pre-trained features for the regression task of DM estimation. The residual connections of the network, or skip connections, enable efficient training of deep architectures by alleviating the problem of vanishing gradients (He et al., 2016b). To repurpose the model for continuous DM prediction, its original classification head, consisting of 1,000 neurons corresponding to ImageNet classes, was replaced with a single linear output neuron. The earlier convolutional layers, which capture general image features such as edges and textures, were kept frozen to retain their pre-trained weights, while the final two convolutional stages (Layers 4 and 5) were unfrozen and trained to extract task-specific patterns from the FRB data. We used the AdamW optimizer with weight decay, which provides an effective and stable form of regularization by directly penalizing large weights, thereby reducing overfitting and preventing the model from overreacting to specific features. We also employed a learning-rate scheduler that reduces the learning rate by a factor of 0.4 every ten training epochs, enabling stable convergence over extended training and allowing further exploration of the model’s learning capacity.
To ensure compatibility with the pre-trained model, the FRB data were pre-processed as follows. Each image was normalised using min–max scaling to the range [0, 1], resized to 224 $\times$ 224 pixels, and replicated across three channels to satisfy the RGB input requirements of ResNet-50. The images were then standardised using the ImageNet mean and standard deviation values ([0.485, 0.456, 0.406] and [0.229, 0.224, 0.225], respectively). To enhance generalisation, a noise matrix was added to the FRB data at each epoch, introducing variability in the noise pattern and emulating the conditions of real FRB observations. The model was trained using the AdamW optimiser (Adam et al., 2014) with a learning rate of $10^{-4}$ and the MAE loss function, which is well suited to continuous regression tasks. Training was performed for 150 epochs, and model performance was monitored using MAE and RMSE metrics.
@@ 4.3 CNN–LSTM Hybrid Model
The recurrent component utilises Long Short-Term Memory (LSTM) units (Hochreiter & Schmidhuber, 1997), which mitigate the vanishing-gradient problem of standard recurrent neural networks by introducing gating mechanisms that control information flow through time. Each LSTM cell maintains two information pathways: the cell state, which stores long-term memory, and the hidden state, which captures short-term dynamics. This dual mechanism enables the network to retain context across the entire 1024 time steps of the input waterfall plots, effectively capturing both transient spectral variations and long-term temporal trends.

### v2

@@ Abstract
pubyear: 2026pagerange: Machine-learning approaches to dispersion measure estimation for fast radio burstsD
@@ 1 Introduction
In this paper, we investigate the use of machine-learning (ML) techniques to improve dedispersion and DM estimation for FRBs. While convolutional neural networks (CNNs; Connor and van Leeuwen 2018) have previously been proposed as substitutes for traditional dedispersion backends, most ML studies to date have focused on the classification or detection of FRBs rather than on quantitative DM estimation. Here, we extend this exploration by developing and comparing three deep-learning architectures, a baseline CNN, a fine-tuned ResNet-50, and a hybrid CNN–LSTM model, designed to recover the intrinsic DM directly from frequency–time dynamic spectra. Section 2 reviews previous ML applications to FRBs, Section 3 describes the generation of synthetic FRB data used for training and validation, and Section 4 outlines the architecture and training procedures for the three models. Our results are presented and discussed in Section 5, and the main conclusions and future directions are summarised in Section 6. Additional details on the simulated data, cross-validation, and test-set performance are provided in the Appendices at the end.
@@ 2 FRBs and Machine Learning
As previously stated, ML techniques have increasingly been applied to FRBs, though most studies have focused on classification rather than quantitative parameter estimation. Existing applications fall into two main categories: distinguishing genuine FRB signals from radio-frequency interference (RFI) or known pulsars (Wagstaff et al., 2016; Akeret et al., 2017; Mosiane et al., 2017; Czech et al., 2018a; Agarwal et al., 2020), and differentiating between repeating and non-repeating sources (Yang et al., 2023; Zhu-Ge et al., 2023; Luo et al., 2023; García et al., 2024; Sharma and Rajpaul, 2024; Qiang et al., 2025; Saliwanchik and Slosar, 2022). Various ML models have been explored in these contexts. For example, Wagstaff et al. (2016) used a Random Forest classifier to enhance FRB detection within the VLBA’s V-FASTR project (Wayth et al., 2011), while Akeret et al. (2017) applied CNNs to discriminate genuine FRBs from RFI and pulsars. Saliwanchik and Slosar (2022) further used neural networks to classify FRBs as repeating or non-repeating.
A notable example is the work of Agarwal et al. (2020), who demonstrated the use of CNNs and transfer learning to improve FRB detection. They fine-tuned pre-trained image-classification networks such as VGG16 (Simonyan and Zisserman, 2015), ResNet50 (He et al., 2016b), and Xception (Chollet, 2017) on a combined dataset of simulated FRBs, real RFI events from the Green Bank Observatory, and pulsar data. Their approach incorporated both frequency–time and DM–time representations to capture spectro-temporal and dispersion-related features, introducing a fusion technique that combined information from the two input domains to enhance classification accuracy.
Interest in ML-driven FRB analysis continues to grow, but the development of more sophisticated models has been limited by the scarcity of publicly available, high-quality data. To overcome this, many studies have generated synthetic data to supplement real observations. For instance, Connor and van Leeuwen (2018) developed a deep CNN trained on dynamic spectra, pulse profiles, and DM–time data using simulated FRBs injected into actual survey noise. They also explored whether CNNs could eventually replace traditional real-time dedispersion backends, but emphasized that low signal-to-noise per pixel remains a fundamental limitation for such applications with current telescopes.
@@ 3.1 Data Simulation
Scattering effects, which both broaden the FRB pulse and slightly delay its peak (Jing et al., 2025), were incorporated by convolving the intrinsic Gaussian profile with a one-sided exponential kernel,
@@
Figure 1: An FRB simulation with $\mathrm{DM}=105.362~\mathrm{pc\,cm^{-3}}$. The background noise is removed to enhance signal visibility while preserving the original sampling time.
@@ 4.1 Baseline Convolutional Neural Network (CNN) Model
The baseline architecture follows the original design of a convolutional neural network (CNN; LeCun et al. 1998) developed by Rajabi (2024), implemented here for DM estimation from frequency–time dynamic spectra. The model comprises seven convolutional layers followed by a sequence of fully connected (dense) layers, as summarised in Table 1. Each convolutional layer applies a set of filters (kernels) to detect local spectro–temporal features and hierarchically capture complex patterns in the input data. To stabilise training and accelerate convergence, batch normalisation (Ioffe and Szegedy, 2015) is applied after each convolutional layer, followed by a Rectified Linear Unit (ReLU) activation (Nair and Hinton, 2010) to introduce non-linearity. Max-pooling operations with $2\times 2$ kernels are used between most convolutional layers (except the third and sixth) to reduce the spatial resolution of feature maps while retaining the most significant features, thereby improving computational efficiency and generalisation.
The convolutional layers progressively increase in depth, starting with 32 filters in the first layer and doubling in number through successive layers, culminating in 2048 filters in the seventh (final) convolutional layer, following the same principle as VGGNet (Visual Geometry Group; Simonyan and Zisserman 2014). This hierarchical structure enables the network to learn increasingly abstract representations, from simple spectral features in early layers to complex temporal correlations in deeper ones. Following the convolutional block, a series of fully connected (FC) layers integrates the extracted features to produce a single scalar output corresponding to the predicted DM. The dense section reads the flattened activations (100,352 elements) from the convolutional block, encodes them into 16,384 outputs, and then halves in size through successive layers until reaching the output neuron. The complete layer-by-layer configuration is summarised in Table 1 (adapted from Rajabi 2024), where the numbers of inputs and outputs ($a{\to}b$) indicate the number of feature maps entering and leaving each convolutional or dense layer.
@@
During hyper-parameter tuning, we explored architectures ranging from four to seven convolutional layers and varied the widths of the dense layers to achieve an optimal balance between model complexity and overfitting. Batch normalisation significantly improved convergence stability by reducing internal covariate shift (Ioffe and Szegedy, 2015), enabling the use of higher learning rates without divergence. Although dropout is a common regularisation technique (Srivastava et al., 2014), tests with dropout probabilities between 0 and 0.2 showed that random unit removal degraded performance, since strong spatial dependencies exist between neighbouring pixels in the input data (Zhang et al., 2018). Consequently, dropout layers were omitted.
@@
The model minimises the MAE loss function, which is well suited to continuous regression tasks such as DM estimation. We used the Adam algorithm (Adam and others, 2014) as the optimiser. Adam adapts learning rates for individual parameters based on mini-batch updates, making it well suited for large datasets in deep learning. Learning rates in the range $10^{-7},10^{-2}]$ were tested, with the smallest value ($10^{-7}$) yielding the smoothest convergence and most accurate results. Larger learning rates led to divergence from the optimal solution, whereas smaller rates allowed for gradual and stable parameter updates.
@@ 4.2 ResNet-50
ResNet-50, a deep convolutional neural network comprising 50 layers (He et al., 2016a), was fine-tuned to adapt its pre-trained features for the regression task of DM estimation. The residual connections of the network, or skip connections, enable efficient training of deep architectures by alleviating the problem of vanishing gradients (He et al., 2016a). To repurpose the model for continuous DM prediction, its original classification head, consisting of 1,000 neurons corresponding to ImageNet classes, was replaced with a single linear output neuron. The earlier convolutional layers, which capture general image features such as edges and textures, were kept frozen to retain their pre-trained weights, while the final two convolutional stages (Layers 4 and 5) were unfrozen and trained to extract task-specific patterns from the FRB data. We used the AdamW optimizer with weight decay, which provides an effective and stable form of regularization by directly penalizing large weights, thereby reducing overfitting and preventing the model from overreacting to specific features. We also employed a learning-rate scheduler that reduces the learning rate by a factor of 0.4 every ten training epochs, enabling stable convergence over extended training and allowing further exploration of the model’s learning capacity.
To ensure compatibility with the pre-trained model, the FRB data were pre-processed as follows. Each image was normalised using min–max scaling to the range [0, 1], resized to 224 $\times$ 224 pixels, and replicated across three channels to satisfy the RGB input requirements of ResNet-50. The images were then standardised using the ImageNet mean and standard deviation values ([0.485, 0.456, 0.406] and [0.229, 0.224, 0.225], respectively). To enhance generalisation, a noise matrix was added to the FRB data at each epoch, introducing variability in the noise pattern and emulating the conditions of real FRB observations. The model was trained using the AdamW optimiser (Adam and others, 2014) with a learning rate of $10^{-4}$ and the MAE loss function, which is well suited to continuous regression tasks. Training was performed for 150 epochs, and model performance was monitored using MAE and RMSE metrics.
@@ 4.3 CNN–LSTM Hybrid Model
The recurrent component utilises Long Short-Term Memory (LSTM) units (Hochreiter and Schmidhuber, 1997), which mitigate the vanishing-gradient problem of standard recurrent neural networks by introducing gating mechanisms that control information flow through time. Each LSTM cell maintains two information pathways: the cell state, which stores long-term memory, and the hidden state, which captures short-term dynamics. This dual mechanism enables the network to retain context across the entire 1024 time steps of the input waterfall plots, effectively capturing both transient spectral variations and long-term temporal trends.